Calculating deposit needed   You put

10,000

in an account earning

6​%.

After

4

​years, you make another deposit into the same account.  

Four

years later​ (that is,

8

years after your original

​$10,000

​deposit), the account balance is

$21,000.

What was the amount of the deposit at the end of year

4​?

The amount of the deposit at the end of year

4

is

​$___.

​(Round to the nearest​ cent.)
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

21000=10000*(1.06)^8+P*(1.06)^4

21000=(10000*1.593848075)+P*(1.06)^4

P=(21000-15938.48075)/1.06^4

=$4009.20(Approx).